OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x – 3y = –3 2x – 6y = –6

OpenStudy (anonymous):

Multiply first equation with 2 and second equation with 5 and then subtract them if you really want the value of y..

OpenStudy (anonymous):

Do you multiply everything ?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

See first one I solve for you: I said multiply first equation by 2, 2(5x - 3y = 3) 10x - 6y = 6 Now do the same for second equation but multiply it with 5..

OpenStudy (anonymous):

Sorry gotta go... @lgbasallote please help @Caramel_823 I have to go now...

OpenStudy (lgbasallote):

so what's left here?

OpenStudy (anonymous):

5(2x-6y)=-6 10x-30y=-30

OpenStudy (lgbasallote):

wait...the system is 5x - 3y = -3 2x - 6y = -6 right?

OpenStudy (anonymous):

Yes so after she multiplied everything how come her last number stayed positive ?

OpenStudy (lgbasallote):

can we start from beginning?

OpenStudy (lgbasallote):

which wouldyou like to cancel? x or y?

OpenStudy (anonymous):

x because I'm trying to find y

OpenStudy (lgbasallote):

so that was why you multiplied by 5...okay let's do that 5x - 3y = -3 10x - 30y = -30 now what would you multiply the first equation so you can cancel out x?

OpenStudy (anonymous):

I think 2 would cancel it out

OpenStudy (lgbasallote):

okay let's try 10x - 6y = -6 10x -30y = -30 so how can we cancel x? what operation do we use?

OpenStudy (anonymous):

multiplication

OpenStudy (lgbasallote):

hmm if you multiply the two equations you'll get a VERY BIG number haha

OpenStudy (lgbasallote):

and the choices are addition or subtraction

OpenStudy (anonymous):

subtraction

OpenStudy (lgbasallote):

you can only use those operation in elimination method remember?

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

ok?

OpenStudy (anonymous):

what would be the next step ?

OpenStudy (lgbasallote):

you said..subtraction..so subtract

OpenStudy (lgbasallote):

subtract the equations 10x - 6y = -6 10x - 30y = -30

OpenStudy (anonymous):

x-24y=-24

OpenStudy (lgbasallote):

hmmm i think you did something wrong in subtraction

OpenStudy (lgbasallote):

first what is 10x - 10x?

OpenStudy (anonymous):

They cancel out , right ?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

now what's -6y - (-30y)?

OpenStudy (anonymous):

24

OpenStudy (lgbasallote):

24y*

OpenStudy (lgbasallote):

what is -6 - (-30)?

OpenStudy (anonymous):

24

OpenStudy (lgbasallote):

therefore what is the difference of 10x - 6y = -6 - 10x - 30y = -30

OpenStudy (anonymous):

24y=24

OpenStudy (lgbasallote):

great..so now solve for y

OpenStudy (anonymous):

y=1

OpenStudy (lgbasallote):

correct

OpenStudy (anonymous):

Thank you !

OpenStudy (lgbasallote):

<tips hat>